OpenStudy (anonymous):

find the lim as x approaches 0 of x squared divided by sin squared x

OpenStudy (irishboy123):

this is a famous limit, just up side down and squared \[\lim\limits_{x \to 0}\frac{x^2}{\sin^2 x} \] do you have any ideas? you cannot use L'Hopital for this one. in fact i don't think you can use calculus at all for this one. sticking my neck out there but makes sense to me! @freckles @Empty @ganeshie8

OpenStudy (freckles):

\[\frac{x^2}{\sin^2(x)}=\frac{x}{\sin(x)} \cdot \frac{x}{\sin(x)} \\ \text{ you should be able to use } \lim_{x \rightarrow 0} \frac{x}{\sin(x)}=1 \]

OpenStudy (anonymous):

isnt it lim x approaches o sin x over x =1

OpenStudy (anonymous):

because my professor wants us to use that theorem but i can't figure out how to do it

OpenStudy (freckles):

the one I mentioned is also true \[\frac{x}{\sin(x)}=\frac{1}{\frac{\sin(x)}{x}} \rightarrow \frac{1}{1} =1 \text{ as } x \rightarrow 0\]

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

for this equation that you just gave me, in order to get the 1 over the sinx divided by x did you divide by 1?

OpenStudy (freckles):

you could divide top and bottom by x but I just know my reciprocals :p example \[x=\frac{1}{\frac{1}{x}} \text{ or } 5=\frac{1}{\frac{1}{5}}\]

OpenStudy (anonymous):

duhhh omg i understand now

OpenStudy (anonymous):

thanks @freckles and @IrishBoy123

OpenStudy (freckles):

np